OpenStudy (anonymous):

Can someone please show me how to solve this problem: A piece of wire 11 2/3 long is divided into two parts such that one part is 2/3 that of the other. find the length of the shorter piece.

OpenStudy (anonymous):

you have a wire of 22/3 long and dive into two parts but, one is 2/3 of the other so: (======================) - wire |---------x-----------|----(2/3)x-| - 22/3 therefore, x + (2/3)x = 22/3 (5/3)x = 22/3 x = (22/3)*(3/5) x = 22/5 = 4 (2/5) The smaller part is then: (22/5)*(2/3) = 44/15 = 2 (14/15)

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

The answer is 4 2/3. Hey.. 11 2/3 is not 35/3?

OpenStudy (anonymous):

ops, yea 11 2/3 is 35/3. Just change (22/3) to (35/3) and you should get the right answer. (x=7)